Bonita, Inc. manufactures ergonomically designed computer furniture. Bonita uses a job order costing system. On November 30,the drop down box's options are: overapplied or underapplied
Answer #1
Predetermined overhead rate = Estimated manufacturing overhead / Estimated machine hours = 4800000 / 960000 5
Total machine hours worked in December = 17000 + 5000 + 24200 + 14000 60200
Total machine hours worked during the year = Machine hours worked till November + Total machine hours worked in December = 842000 + 60200 902200
Total actual manufacturing overhead incurred during the year = Manufacturing overhead incurred till November + Manufacturing overhead incurred in December = 4294500 + 395000 4689500
Total manufacturing overhead applied during the year = Total machine hours worked during the year * Predetermined overhead rate = 902200 * 5 4511000
As the manufacturing overhead applied is less than the actual manufacturing overhead, the overhead is underapplied by 178500 [ i.e. 4689500 - 4511000 ]
Overhead underapplied by $178,500